Please help me with the correct explanation and correct answer. ThanksElectroplating It takes 207.8 s for a 1.07-A current to plate 0.1368 g of a metallic element from a solution containing M2+ i
Answer #1

let us calculate the charge passed:

time, t = 2.078*10^2s

Q = I*t

= 1.07A * 2.078*10^2s

= 2.223*10^2 C

Electrolysis equation is:

M2+ + 2e- ------> M

1 mol of M requires 2 mol of electron

1 mol of electron = 96485 C

So,1 mol of M requires 192970 C

mol of M plated = 2.223*10^2/192970 = 1.152*10^-3 mol

use:

number of mol = mass / molar mass

0.0012 = 0.1368/molar mass

molar mass = 118.7 g/mol

This is molar mass of tin

Answer: Sn